# Neutral residues: Revisiting adapters for model extension

### v1

@@ 4 Experiments
In this section, we report experimental results to validate the performance of neutral residues to adapt a neural network to a new domain. We consider the case of adding or improving the multilingual ability of a large language model. More precisely, we start from an English-only model (or a model that has only seen a small amount of non-English data) and finetune it on French data.
@@ 4.1 Datasets and evaluation protocols
For the French finetuning dataset, we use a dataset extracted from CommonCrawl, which was pre-processed with the following steps. First, the text content was extracted from HTML using the resiliparse package. Then, we performed language identification with fastText to keep only French data. Next, we performed deduplication at the paragraph level, by computing a hash of each paragraph. Finally, we filtered document using a fastText linear classifier, which was trained to discriminate random pages from CommonCrawl vs. high-quality documents such as Wikipedia articles, textbooks, news articles or scientific articles.
For the English domain, we restricted ourselves to text from Wikipedia. The motivation for this is to use a finetuning dataset that is close, but yet different, from the training dataset of the backbone model. Indeed, as discussed earlier, in many cases one does not have access to the original dataset that was used to train the backbone model, and we want to demonstrate that our method is robust even in the case where exact data from the original domain is not available.
@@
We consider two main ways to evaluate the performance of finetuning. First, we evaluate the perplexity of the model on held-out sets, both in English and in French, to measure forgetting and learning. For English, we consider text from a domain that does not correspond to the finetuning data, and use the PubMed subset from ThePile (Gao et al., 2020). The goal here is to evaluate how robust our method is to forgetting, especially in the case where we do not have access to the same training distribution as the original model. For French and German, we consider text from the same distribution as the finetuning dataset.
Second, we consider standard academic benchmarks used to evaluate large language models, such as question answering or Cloze-style problems. We use the following datasets: ARC challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020), CommonSense QA (Talmor et al., 2018) and Belebele (Bandarkar et al., 2023). For French, we use the translated datasets from Dac Lai et al. (2023) and Sakai et al. (2024). For all datasets, we score each answer independently as the continuation of the question, and predict the most probable one. We use normalization, following the recommendation from Gu et al. (2024). We perform 5-shot evaluation for ARC challenge, CSQA and MMLU, and 0-shot evaluation for HellaSwag and Belebele.
Backbone models.
We consider three models that we use as backbone. The two first models were trained internally on 2.4T tokens, and have 1B parameters. The main difference between these two models is that the first model, denoted Transf-EN, was trained on English data only. The second model Trans-ML is MultiLingual: it was trained on a mix of English, French, German, Spanish, Italian and Portuguese data. The non-English data represents roughly 12.5% of the tokens. The last model is Gemma-2B (Gemma et al., 2024). It also includes a small amount of multi-lingual data during pre-training. We do not have access to the training dataset of Gemma, making it a realistic test case for our approach. Appendix A gives more details about these models.
For finetuning with new knowledge, we train during 100000 steps with a batch size of 64 for Trans-ML and Trans-EN, and we train Gemma-2B during 150000 steps with a batch size of 8.
Baseline extension strategies.
In our experiments, we use the following method as baselines:
  • LoRA: unlike Hu et al. (2022), we add all the additional parameters in the FFN layers, as in our preliminary experiments this choice was significantly better for knowledge extension.
  • Adapters: we use the vanilla adapters as Houlsby et al. (2019) with near-identical initialization. We use parallel instead of serial blocks to have a more directly comparable baseline.
Hyperparameters.
Except mentioned otherwise, LoRA and both adapters use 20% of extra learnable weights. We use a learning rate of $5\cdot 10^{-5}$ by default, except for finetuning since the trade-off between learning and not forgetting is better with a learning rate of $2\cdot 10^{-4}$, see ablations in Section 4.4. The hyperparameters $\alpha$ and $p$ have been selected based on perplexity results on the validation set. We provide other training hyperparameters in Section A.
4.2 Preliminary Analysis
Training with mixed data distribution
Table 1: Mixture of distributions: Impact of the rate $p$ of data that follows the pretraining distribution when training a vanilla adapter. 10% of initial data distribution is a good compromise between learning (Fr ppl) and not forgetting (En ppl).
@@ 4.2 Preliminary Analysis
Table 2: Training with new data only or mixed data. Starting from an English pretrained model, we measure the tradeoff in between learning and not forgetting if we only post-train with French or if we keep $p=10\%$ of English. We consider full finetuning, LoRA and vanilla adapters baseline, and our neutral residues. See Table 2 for results of the pretrained model. We report the perplexity on the validation sets at the end of training.
@@
Table 2 analyzes the impact of the proportion $p$ of data from the pretraining distribution used to extend the model, measured for vanilla adapters. As one can see, $p=0.1$ offers a good trade-off: the forgetting is not significantly higher than with $p=0.5$, while the model is almost as good in French as when we train with English only ($p=0$). Therefore, when training with mixed data, we set $p=0.1$ in all subsequent experiments.
Table 2 gives the trade-off between learning and not forgetting for all baselines and our method, when training either without or with mixed training distributions. As one can see, the mixed training significantly reduces the forgetting for all the methods without hindering too much the learning.
Table 3: Impact of the number of parameters. We vary the extra capacity allocated to new learnable weights in our method. As to be expected, the capacity of our adapters to perform well in the new language strongly depends on the number of weights that we add. Interestingly, the impact of extra weights on forgetting is almost neutral. Extra Perplexity Tasks avg. weights En Fr En Fr 1+5% 0.667 0.849 46.8 39.5 +10% 0.667 0.819 46.2 40.8 +20% 0.668 0.793 46.5 41.2 +50% 0.669 0.761 46.7 42.3
@@ 4.3 Main results
Table 4 reports our main detailed results when we want to improve the performance in French of a pretrained model. Table 5 provides the counterpart when the new language to be added to the pretrained model is German, which overall leads to the same conclusions. As one can see, our proposal is overall offering the best trade-off between learning and not forgetting compared to all other techniques: It is first in terms of the sum of both metrics (English+French or English+German).
Impact of the initial model.
We observe that the results are consistent for all models. This indicates that even if we do not know exactly the data employed to produce the pretrained model, our neutral residues is beneficial. Note, this is also evidenced by Table 2, where neutral residues perform significantly better than other techniques even when we do not use any pretraining data. If we compare Transf-EN with Trans-ML, as expected the pretrained model Trans-ML obtains much better performance on French (resp. German). However, even the Trans-ML model, which has seen some multilingual at training time, significantly benefits from our posttraining method.
Table 4: Main results: comparison of neutral residues (ours) versus four baselines for three pretrained models. All methods except finetuning use 20% of trainable parameters compared with the backbone models. We use a proportion $p=0.1$ of French for the training. We report the average performance across all tasks both on English and French. These metrics reflect the overall performance with respect to not forgetting and learning, respectively. In bold we report the two best models.
@@
Transf-EN Backbone 43.0 60.3 40.5 56.5 34.5 34.6 32.8 25.9 37.4 27.4 47.0 31.6
finetuning 37.9 45.5 34.8 45.5 31.0 42.2 48.6 34.5 57.0 31.7 39.0 42.8
LoRA 42.0 53.1 38.1 52.3 33.2 42.4 46.2 33.0 51.5 30.7 43.7 40.8
Adapters 42.1 56.9 38.7 53.6 33.4 43.2 46.2 31.7 50.0 30.9 45.0 40.4
ours 43.4 59.5 39.9 55.4 34.3 43.6 48.5 32.2 50.4 31.4 46.5 41.2
Transf-ML Backbone 41.2 59.4 40.6 57.4 34.4 42.6 48.9 34.4 47.6 31.1 46.6 40.9
finetuning 40.1 52.3 37.2 54.3 32.0 43.8 52.7 40.4 57.8 32.6 43.2 45.5
LoRA 40.0 57.1 39.8 54.1 33.5 44.2 51.1 37.4 53.1 31.7 44.9 43.5
Adapters 41.2 58.0 39.0 55.1 34.6 41.8 50.0 38.0 55.4 31.4 45.6 43.3
ours 42.1 59.3 40.6 55.3 34.2 42.9 51.1 38.1 53.7 32.1 46.3 43.6
Gemma-2B Backbone 46.3 69.7 47.0 63.0 39.9 44.3 50.3 36.8 43.8 32.5 53.2 41.5
finetuning 43.7 55.2 38.8 52.6 34.3 45.2 55.3 38.4 56.4 33.7 44.9 45.8
LoRA 48.2 66.7 45.0 60.0 37.9 46.7 55.5 40.5 55.6 35.9 51.6 46.8
Adapters 43.7 66.0 44.3 40.6 38.3 43.7 53.3 34.6 52.4 34.5 46.6 43.7
ours 47.4 69.1 47.6 63.1 40.0 47.7 55.3 41.0 54.5 34.7 53.4 46.6
Table 5: Learning German: we use the same setting as in Table 4 except that we learn German.
Forgetting: English Learning: German Task avg.
Model Method

### v3

@@ 4 Experiments
In this section, we report experimental results to validate the performance of neutral residues to adapt a neural network to a new domain. We consider the case of adding or improving the multilingual ability of a large language model. More precisely, we start from an English-only model (or a model that has only seen a small amount of non-English data) and finetune it on another target language. Section F covers the case of multiple languages for a multilingual model.
@@ 4.1 Datasets and evaluation protocols
For the multi-lingual finetuning datasets, we use data extracted from CommonCrawl, and process it with the following steps. First, the text content is extracted from HTML using the resiliparse package. [^2] https://resiliparse.chatnoir.eu Then, we perform language identification with fastText [^3] https://fasttext.cc to keep data in the target language only. Next, we perform deduplication at the paragraph level, by computing a hash of each paragraph. Finally, we filter document using a fastText linear classifier, which was trained to discriminate random pages from CommonCrawl vs. high-quality documents such as Wikipedia articles, textbooks, news articles or scientific articles.
For the English domain, we restricted ourselves to text from Wikipedia. The motivation for this is to use a finetuning dataset that is related, but different, from the training dataset of the backbone model. Indeed, as discussed earlier, in many cases one does not have access to the original dataset that was used to train the backbone model, and we want to demonstrate that our method is robust even in the case where exact data from the original distribution is not available.
@@
We evaluate the performance of finetuning with two steps. First, we evaluate the perplexity of the model on held-out sets, in English and in the target language, to measure forgetting and learning. For English, we use text from a domain that does not correspond to the finetuning data: the PubMed subset from ThePile (Gao et al., 2020). The goal here is to evaluate how robust our method is to forgetting, especially in the case where we do not have access to the same training distribution as the original model. For target languages, we consider text from the same distribution as the finetuning dataset.
Table 1: Mixture of distributions: impact of the rate $p$ of data that roughly approximate the pretraining distribution when learning a vanilla adapter. 10% of initial data distribution is a good compromise between learning (Fr) and not forgetting (En). We report perplexity (lower is better).
@@
Second, we consider standard academic benchmarks used to evaluate large language models, such as question answering or Cloze-style problems. We use the following datasets: ARC challenge (Clark et al., 2018), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020), CommonSense QA (Talmor et al., 2018) and Belebele (Bandarkar et al., 2023). For target languages, we use the translated datasets from Dac Lai et al. (2023) and Sakai et al. (2024), when they exist. For all datasets, we score each answer independently as the continuation of the question, and predict the most probable one. We use normalization, following the recommendation from Gu et al. (2024). We perform 5-shot evaluation for ARC challenge, CSQA and MMLU, and 0-shot evaluation for HellaSwag and Belebele.
Backbone models.
We consider two models that we use as backbone. The first, called EN-LM-1B, was trained internally on 2.4T tokens of English data only, and has 1B parameters. The pretraining data consists mostly of web pages filtered from Common Crawl, as described previously. It also includes a small amount of data from curated sources such as Wikimedia, Stack Exchange and scientific articles. The second model is Gemma-2B (Gemma et al., 2024) which was pretrained on a small amount of multi-lingual data. We do not have access to Gemma’s training data, making it a realistic test case for our approach. Appendix B gives more details about these models. We use EN-LM-1B for all the experiments, except for the main results that are reported in Table 3. We chose to use EN-LM-1B for our exploratory experiments and ablations as it is smaller, hence leading to faster experiments.
Baseline extension strategies.
In our experiments, we use the following methods as baselines:
  • LoRA: unlike Hu et al. (2022), we add all the additional parameters in the FFN layers, as in our preliminary experiments this choice was significantly better for model extension.
  • Adapters: we use the vanilla adapters as Houlsby et al. (2019). We use parallel instead of serial blocks to have a more directly comparable baseline. We use the standard adapters’ initialization, ie. zeros for the output matrix and He’s initialization for the other weights.
Table 2: Training with new data only or mixed data. Starting from an English pretrained model, we measure the tradeoff between learning and forgetting if we only post-train with target data or if we keep $p=10\%$ of original data. See Table 1 for results of the pretrained model. We report the perplexity on the validation sets at the end of training.
@@
Hyperparameters.
Except mentioned otherwise, LoRA, adapters, and neutral residues use 20% of extra learnable weights. For each method, we selected the learning rate that leads to the best trade-off between learning and forgetting: $5\cdot 10^{-5}$ for finetuning and LoRA, $2\cdot 10^{-4}$ for adapters and neutral residues. The hyperparameter $\alpha$ governing the strength of the sparsity loss is set by default to 0.010.010.01 for neutral residues. When training on the new data, we train during 100,000 steps with a batch size of 64 sequences of length 4,096 for both EN-LM-1B and Gemma-2B. We provide other training hyperparameters in Section B.
Table 3: Main results: comparison of neutral residues (ours) versus four baselines for the Gemma-2B model. All methods except finetuning use 20% of trainable parameters compared with the backbone models. We use a proportion $p=0.1$ of English data for the training. We report the average performance across all tasks both on English and target languages. These metrics reflect the overall performance with respect to not forgetting and learning, respectively. We report the best method in bold, and the second best is underlined.
Forgetting: English Learning: Target Task avg.
Target Method
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
BeleBele (0)
HellaS (0)
Arc C (5)
CSQA (5)
MMLU (5)
Forgetting
Learning
French Backbone 46.3 69.7 47.0 63.0 40.2 44.3 50.3 40.9 53.3 33.9 53.2 44.6
Finetuning 45.7 62.7 45.9 57.3 37.2 44.9 60.8 43.9 63.8 35.9 49.8 49.9
LoRA 46.1 65.7 45.2 60.5 38.4 46.3 57.1 41.7 44.2 35.1 51.2 44.9
Adapters 48.1 67.8 47.5 61.4 39.3 44.6 57.4 35.2 56.9 36.0 52.8 46.0
Ours 48.2 68.5 47.3 63.3 39.2 46.8 57.9 39.7 59.8 36.7 53.3 48.2
Danish Backbone 46.3 69.7 47.0 63.0 40.2 41.6 43.1 35.6 n/a 32.3 53.2 38.1
Finetuning 46.3 62.3 37.9 55.1 34.5 41.3 57.9 39.2 n/a 32.8 47.2 42.8
LoRA 45.2 64.7 44.7 52.4 35.3 42.2 55.3 37.4 n/a 32.9 48.5 41.9
Adapters 46.9 66.8 45.9 57.1 37.9 42.3 53.7 39.0 n/a 34.2 50.9 42.3
Ours 45.4 69.4 46.6 60.4 38.1 42.4 55.8 38.7 n/a 34.7 52.0 42.9
Hungarian Backbone 46.3 69.7 47.0 63.0 40.2 34.3 36.0 31.1 n/a 31.0 53.2 33.1
Finetuning 44.6 60.3 40.8 48.5 35.2 37.1 48.1 36.1 n/a 32.5 45.9 38.5
LoRA 45.0 63.6 43.3 54.6 36.3 39.0 44.9 35.4 n/a 32.7 48.6 38.0
Adapters 48.0 66.5 48.1 59.1 38.1 39.6 44.8 36.6 n/a 32.7 52.0 38.4
Ours 46.1 69.2 48.5 60.6 38.5 38.2 46.1 38.1 n/a 32.8 52.6 38.8
Slovak Backbone 46.3 69.7 47.0 63.0 40.2 38.1 38.8 32.8 n/a 31.3 53.2 35.2
Finetuning 43.2 60.0 40.0 54.3 34.3 41.3 47.8 35.2 n/a 32.4 46.4 39.2
LoRA 44.8 63.8 41.7 54.9 36.0 41.4 46.7 32.9 n/a 32.3 48.2 38.4
Adapters 45.2 62.5 38.0 55.9 37.4 42.3 40.9 34.9 n/a 32.7 47.8 37.7
Ours 47.9 68.0 42.8 60.6 38.1 40.8 46.6 34.7 n/a 32.2 51.5 38.6
4.2 Preliminary Analysis
Training with mixed data distribution
Table 1 analyzes the impact of the ratio $p$ of data approximating the pretraining distribution used to extend the model with vanilla adapters. We note that $p=0.1$ offers a good trade-off: the forgetting is not significantly higher than with $p=0.5$, while the model is almost as good in French as when we train with French only ($p=0$). Thus, when training with mixed data, we set $p=0.1$ in all subsequent experiments.
Table 2 gives the trade-off between learning and not forgetting for all baselines and our method, when training either without or with mixed training distributions. As one can see, the mixed training significantly reduces the forgetting for all the methods without hindering too much the learning. We report more extensive results in Table 10 of the appendix, including on downstream tasks, when evaluating the impact of including some data related to the original distribution.
Table 4: Ablation: Gating and local loss. We report the average across tasks; see Table 16 for details.
Gate Gate loss Perplexity $\downarrow$ Tasks $\uparrow$
$\ell_{1}$ CE En Fr En Fr
(1) $\varnothing$ 0.687 0.801 45.3 42.4
(2) Sigmoid 0.677 0.800 45.2 42.6
(3) 0.676 0.800 45.3 41.9
(4) ReLU 0.673 0.791 46.4 42.8
(5) 0.674 0.791 47.1 43.6
(6) Adapter baseline 0.686 0.812 45.1 41.3
(7) Backbone baseline 0.663 1.175 47.0 32.9
@@ 4.3 Main results
We report our main results in Table 3, where we extend the capabilities of the Gemma 2B model on four languages: Danish, French, Hungarian and Slovak. First, we note that our method, neutral residues, outperforms the LoRA and adapters baselines on all four languages. In average, it is always among the two best methods: it is only second to the backbone on forgetting and to finetuning on learning. However, neutral residues offers a significantly better trade-off than both these methods. Second, we observe that adapters tend to perform better than LoRA on the task of extending Gemma 2B to a target language. This is a confirmation that for tasks requiring significant new knowledge, it is helpful to explicitly add capacity to the base model in the form of additional parameters. This observation is confirmed when varying the number of additional parameters: we observe improvements when increasing the extra capacity from 5% to 50% (we report these results in Table 9 of the Appendix).
Table 5: Ablation: initialization. We report the average across tasks; see Table 17 for details.
Gate Our Perplexity $\downarrow$ Tasks avg. $\uparrow$
init. En Fr En Fr
Adapters w/ $\ell_{1}$-norm 0.687 0.812 45.5 41.6
0.687 0.801 45.3 42.4
Adapters w/ sigmoid 0.668 0.810 45.6 40.8
0.677 0.800 45.2 42.6
Neutral Residues 0.673 0.818 46.5 41.6
0.674 0.791 47.1 43.6
Table 6: Ablation: $\boldsymbol{\alpha}$ governs the strength of the $\ell_{1}$ local loss. See Table 18 for detailed results.
Perplexity $\downarrow$ Tasks avg. $\uparrow$
$\alpha$ En Fr En Fr
0 0.673 0.791 46.4 42.8
0.01 0.674 0.791 47.1 43.6
0.1 0.672 0.791 46.1 42.9
1 0.674 0.791 46.7 42.3
10 0.672 0.791 46.6 43.2
100 0.667 0.791 46.3 43.0
Table 7: Trade-offs with different learning rates. This hyperparameter allows us to favor learning vs. forgetting or vice-versa. These results are the ones reported in Figure 1. See Table 19 for detailed results.
Perplexity $\downarrow$ Tasks avg. $\uparrow$
Method LR En Fr En Fr
Backbone 0 0.663 1.175 47.0 32.9
$2\cdot 10^{-5}$ 0.678 0.963 45.2 36.3
Fine- $5\cdot 10^{-5}$ 0.705 0.840 44.4 41.1
tuning $2\cdot 10^{-4}$ 0.811 0.758 38.9 43.6
$5\cdot 10^{-4}$ 0.874 0.757 36.0 42.5
LoRA $2\cdot 10^{-5}$ 0.716 0.842 44.1 40.5
$5\cdot 10^{-5}$ 0.730 0.819 43.7 41.2
$2\cdot 10^{-4}$ 0.745 0.802 41.9 41.7
$5\cdot 10^{-4}$ 0.755 0.806 41.6 42.6
$2\cdot 10^{-5}$ 0.686 0.830 45.1 40.8
Vanilla $5\cdot 10^{-5}$ 0.686 0.812 45.1 41.3
adapters $2\cdot 10^{-4}$ 0.689 0.797 46.0 42.4
$5\cdot 10^{-4}$ 0.688 0.796 45.4 43.5
$2\cdot 10^{-5}$ 0.673 0.799 46.4 42.9
Neutral $5\cdot 10^{-5}$ 0.674 0.791 47.1 43.6
residues $2\cdot 10^{-4}$ 0.673 0.787 46.4 42.5
$5\cdot 10^{-4}$ 0.675 0.789 46.6 43.0
4.4 Ablations
Ablation of the gating and local loss.
In Table 4, we compare adapters using different gating mechanisms, and different loss functions to learn the gating. First, comparing line (5) to line (1), we observe that regularizing the output of the adapters alone, without adding a block gating mechanism does not work well. We hypothesize that this is mainly due to the weights of the adapter to become co-linear, as discussed in spectral analysis of Section 3. In particular, this prevent the adapter to learn the new domain well. Next, we observe that the combination of the ReLU activation with $\ell_{1}$-norm outperforms the sigmoid activation with cross entropy loss (line (5) vs. (2)). Finally, comparing lines (4) and (5), we note that adding supervision with the $\ell_{1}$ regularization helps learning adapters, even though the ReLU gating alone already leads to strong results.
Initialization ablation.
In Table 5, we report the performance of various adapters, with and without the low-variance initialization described in Section 3. Overall, we note that this initialization improves the results of gated adapters. More specifically, compared to He’s initialization, it leads to better learning of the new target domain.
Learning rate ablation.
Table 7 reports the trade-offs achievable by different techniques when varying the learning rate. The results of our method are stable across a large range of learning rates (one order of magnitude from $5\cdot 10^{-5}$ and $5\cdot 10^{-4}$), in sharp contrast with finetuning and LoRA, which are highly sensitive to this parameter.
Coefficient $\boldsymbol{\alpha}$ ablation.
Table 6 shows that even without the sparsity constraint, ReLU gating enables a good balance between learning and forgetting. Introducing this constraint further improves the trade-off, with low values ($\alpha=0.01$) yielding the best results. However, a good trade-off is maintained across a wide range of strengths.
5 Conclusion
This paper has explored the feasibility and effectiveness of extending existing foundation models to incorporate new capabilities or knowledge without the need for resource-intensive retraining from scratch. By building upon the concept of adapters, which add new parameters to a pretrained model, we have demonstrated that it is possible to extend a model without compromising its original knowledge, thereby offering a more sustainable approach than retraining from scratch. Our study focused on the use-case of adding a new language to a pretrained model and evaluated the performance of the extended model on both training criteria and downstream tasks. Our findings highlight several critical factors that contribute to the successful extension of a model while mitigating the issue of catastrophic forgetting. These factors include the strategic combination of mixed training data, an adapter gating mechanism coupled with a local loss, and the importance of near-identical initialization. Our method, called neutral residues, obtains state-of-the-art results on extending multilingual capabilities of Gemma-2B.
Impact Statement
The goal of our work is to advance the efficiency of large langage models (LLMs) training, by improving methods to extend an existing model to a domain requiring a large amount of new knowledge. This has the potential to improve the sustainability and the accessibility of these systems, across various applications and domains. In particular, this has potential environmental implications, as it could reduce the computational resources needed to develop a model for a new domain. Instead of training the model from scratch, one could start from a strong existing base model and use our method to extend it to the target domain. Our method can also improve the accessibility of these systems, by allowing organizations and researchers to personalize existing models to their need without requiring a large amount of computational resources.
In general, the widespread application of LLMs has many other potential social consequences, including risks. It requires careful considerations of ethical issues around fairness, transparency and potential misuse. As our research do not directly address any of these, we do not feel any of these risks must be specifically highlighted here. We thus refer the reader to previous work on the matter, such as Bommasani et al. (2021) or Bender et al. (2021).
Appendix A Architecture details and Hyper-parameter settings
We consider a vanilla transformer EN-LM-1B, which was trained internally on 2.4T tokens of English data only, and has 1B parameters. We also considered the Gemma 2B model (Gemma et al., 2024) and Gemma2 2B model (Team et al., 2024), for which we do not know the distribution. For these models, we refer to their paper for training hyper-parameters.
Table 8: Hyper-parameters details
EN-LM-1B Gemma 2B Gemma2 2B
Number of layers $L$ 16 18 18
Working dimensionality 2048 2048 2048
Number of heads 128 256 256
Dimension FFN latent space 5632 16384 16384
Activation type Gated SILU Gated GELU Gated GELU
Normalization RMS pre-norm RMS pre-norm RMS pre & post norm
Group query
Pretraining context length 4096 8192 8192
Hyper-parameters for training the extended model
Batch size 64 64 64
Context length 4096 4096 4096
#Training steps 100,000 100,000 100,000
AdamW: $\beta_{1}$ 0.9 0.9 0.9
AdamW: $\beta_{2}$ 0.95 0.95 0.95
Appendix B Impact of the number of parameters
We vary the extra capacity allocated to new learnable weights in LoRA, adapters and neutral residues (ours). The steady performance improvement in the new language highlights the importance of expanding the base model’s capacity for tasks requiring substantial new knowledge. Interestingly, the impact of extra weights on forgetting is almost neutral. We trained EN-LM-1B with a learning rate of $2\cdot 10^{-4}$ for finetuning and adapters, and $5\cdot 10^{-5}$ for LoRA and ours with other hyperparameters detailed in Table 8. For each extra weight size, we report the best method in bold.
Table 9: Impact of the number of parameters.
@@ Appendix B Impact of the number of parameters
Backbone 43.0 60.3 40.5 56.5 34.5 34.6 32.8 26.7 42.3 28.0 47.0 32.9
Finetuning 37.9 45.5 34.8 45.5 30.9 42.2 48.6 35.3 60.3 31.5 38.9 43.6
+5% LoRA 42.0 54.1 36.1 50.1 32.8 41.7 42.3 32.4 50.0 30.4 43.0 39.4
Adapters 42.2 57.9 40.4 55.1 33.7 44.0 44.1 34.2 49.5 30.5 45.9 40.5
ours 42.9 59.3 40.1 55.8 33.8 42.2 44.1 33.6 52.1 30.3 46.4 40.5
+10% LoRA 41.7 53.6 38.1 51.1 32.8 42.4 44.2 32.2 51.8 30.3 43.5 40.2
Adapters 42.9 57.5 38.7 53.3 33.6 43.2 46.4 33.5 53.4 30.7 45.2 41.5
ours 44.4 58.9 42.2 56.8 34.2 44.3 46.5 34.1 56.5 31.0 47.3 42.5
+20% LoRA 42.0 53.1 38.1 52.3 32.8 42.4 46.2 32.7 54.2 30.7 43.7 41.2
Adapters 43.1 57.7 40.3 54.1 34.5 44.3 48.8 34.8 52.7 31.2 46.0 42.4
ours 43.9 59.3 41.6 56.8 34.1 45.1 48.8 35.7 56.9 31.6 47.1 43.6
+50% LoRA 41.8 52.0 36.1 51.0 32.7 44.3 47.9 35.6 56.2 30.8 42.7 43.0
Adapters 43.3 57.6 39.4 53.3 33.5 45.8 50.4 36.5 58.0 31.9 45.4 44.5
ours 43.1 59.2 40.3 56.3 34.2 43.8 50.9 35.2 59.3 32.1 46.6 44.3
Appendix C Impact of mixed data distribution.
Table 10 shows the importance of training with data approximating the pretraining distribution of the model. The mixed training significantly reduces the forgetting for all the methods without hindering too much the learning. We trained EN-LM-1B with a learning rate of $2\cdot 10^{-4}$ for finetuning and adapters, and $5\cdot 10^{-5}$ for LoRA and ours with other hyperparameters detailed in Table 8. We report the best method in bold, and the second best is underlined
Table 10: Impact of mixed data distribution.
Forgetting: English Learning: French Task avg.
Ratio Method